OpenStudy (anonymous):

Write a equation the sum of two number is 27. The large number is 3 more than the smaller number . What are the two numbers ?

OpenStudy (anonymous):

x+y=27 y+3=x (y+3)+y=27 2y+3=27 Now you solve!